Use an inverse matrix to solve the system of equations, if possible. 5x + 4y + z = -73 3.x - 6y + 3z = 45 -4x+8y – z= -33

Choose one:

a. (-5, -10, -8)

b. (-10, -8, 9)

c. (-5, 3, 9)

d. No Solution
Answer #1

-4x+8y-z=-38 - + 7 =-73 - 0 64 +37 45 > x-2y+z = 15 - ① 50+44 3a A = 1 From this 3 equation we get, -73 45 - -33 the inverse